A firm has a weighted average cost of capital of 10 percent. The firm's aftertax cost of debt is 5.8 percent and the cost of equity is 11.8 percent. The tax rate is 34 percent. What is the firm's debt-equity ratio?
1
Expert's answer
2012-02-28T07:38:56-0500
The debt-to-equity ratio (D/E) is a financial ratio indicating the relative
proportion of shareholders' equity and debt used to finance a company's assets:
D/E = Debt(liabilities)/equity.

WACC = (Debt / (Debt + Equity)) * Cost of
Debt + (Equity / (Debt + Equity)) * Cost of Equity

Cost of Debt =
After-tax Cost of Debt / (1- Tax Rate) = 0.058/(1-0.34) = 0.088 or
8.8%

D/E = 1.5.